If m^2x+1=m^9 what is the value of x?

If possible, enter your answer as an integer, like this: 42

If your answer must be rounded, enter it as a decimal rounded to one place, like this: 42.5


(i got 11, but it's not right)



1
Expert's answer
2021-06-15T09:29:14-0400

Let us solve the equation "m^{2x+1}=m^9." It follows that "2x+1=9," and hence "2x=8." We conclude that the solution of the equation is "x=4."